# Measuring Ransomware Lateral Movement Susceptibility via Privilege-Weighted Adjacency Matrix Exponentiation Satyam Tyagi1 Ganesh Murugesan1

### v1

@@ 6.  Examples and Evidence
In this section, we provide a practical demonstration of our model. We begin with a simple, illustrative example to clarify the computational steps. We then present a more complex scenario to provide evidence of the model’s utility in identifying non-obvious security risks that are missed by standard reachability analysis.
6.1 Illustrative Example
Consider a simple network of three assets: a Web Server (Asset 1), an Application Server (Asset 2), and a Database (Asset 3). For this example, the Adjacency Matrix ($A$) representing the reachable services, and the corresponding Pivot Potential Matrix ($W$) derived from those services, are defined as:
$A=\begin{bmatrix}0&1&0\\ 1&0&1\\ 0&0&0\end{bmatrix}\quad W=\begin{bmatrix}0.0&0.2&0.0\\ 0.9&0.0&0.3\\ 0.0&0.0&0.0\end{bmatrix}$
Using the iterative algorithm defined in Section 3, we compute the 2-hop compromise probability matrix, $P_{2}$:
  • First, the 1-hop matrix $P_{1}$ is initialized from $A$.
  • Then, we compute the 2-hop result: $P_{2}=P_{1}\lor_{p}(P_{1}\otimes W)$.
This yields the final probability matrix:
$P_{2}=\begin{bmatrix}1.0&1.0&0.3\\ 1.0&1.0&1.0\\ 0.0&0.0&1.0\end{bmatrix}$
From this matrix, we calculate the final metrics for $K=2$:
  • LMS2: The average of the 6 off-diagonal elements is $(1.0+0.3+1.0+1.0+0.0+0.0)/6=0.55$.
  • BRE2: The average of all 9 elements is $(6.3)/9\times 100\%=70\%$.
This simple example demonstrates how the high pivot potential of the connection from Asset 2 to 1 creates a high probability of compromise across the network, resulting in a high Blast-Radius Estimate.
@@ 6.2 Evidence of Practical Utility
We now provide evidence of the model’s value in a more complex scenario where the goal is to prioritize security controls. Consider a network with four assets: a User Workstation (Asset 1), a Web Server (Asset 2), a critical Database (Asset 3), and a Domain Controller (Asset 4).
The reachable services are:
  • Workstation (1) $\to$ Web Server (2) via a low-privilege web port, $\pi=0.1$.
  • Web Server (2) $\to$ Database (3) via a standard SQL port, $\pi=0.2$.
  • Workstation (1) $\to$ Domain Controller (4) via an RDP port left open for remote admin access, $\pi=0.9$.
Analysis with the Unweighted Method: A standard reachability analysis (using the Boolean method) would identify two lateral movement paths originating from the Workstation: one to the Database (length 2) and one to the Domain Controller (length 1). Both are considered ”reachable.” A security team might mistakenly prioritize the longer path involving the Web Server, or treat both paths as equally important, failing to see the true concentration of exposure.
Analysis with Our Probabilistic Model: Our model calculates the compromise probabilities. The 2-hop compromise matrix $P_{2}$ would show:
  • The probability of compromising the Database from the Workstation, $(P_{2})_{13}$, would be low. The probability of the two-hop path is based on the composition of low-privilege pivots (approximately $0.1\times 0.2=0.02$).
  • The probability of compromising the Domain Controller, $(P_{2})_{14}$, would be very high, dominated by the single high-privilege RDP hop (0.9).
The BRE from the Workstation would be heavily weighted by the near-certain compromise of the Domain Controller.
@@ 6.3 Illustrative Field Example: Service-Class Contrast in an Enterprise Snapshot
We complement our formal results with an illustrative observation from a production environment described by Tyagi [2025]. The analysis compares adjacency induced by two services: RDP (TCP/3389), an interactive remote-admin channel, and Epic Cache (TCP/6966), an application-specific service. The induced RDP adjacency was reported as nearly complete (most node pairs reachable directly), whereas the Epic Cache adjacency was nearly empty (few pairs reachable). Under our framework, this service-class asymmetry implies substantially higher step-success and union mass for RDP-driven paths, elevating both LMSK (off-diagonal mean) and BREK (mean including the diagonal) relative to app-only services.
The same field note underscores that adversaries frequently achieve lateral movement using valid accounts over remote services (RDP/SSH/SMB/WinRM) without relying on unpatched CVEs, aligning with ATT&CK’s “Remote Services” technique and our CVE-agnostic design. We emphasize that this is a single-environment qualitative example (not a statistical study). Nevertheless, the observation is consistent with our model’s predictions and with practitioner guidance prioritizing constraints on interactive remote services. [Tyagi, 2025; MITRE, 2025b, f, c]
Conclusion: This scenario provides clear evidence of the model’s value. While the unweighted method presents a noisy picture of undifferentiated risks, our probabilistic model immediately and quantitatively identifies the single RDP connection as the most significant threat to the network. It correctly filters the noise from the low-probability path and pinpoints the connection that must be prioritized for remediation, demonstrating its practical utility in guiding security decisions.

### v2

@@ 7.1 Evidence of Practical Utility
We reuse the running 4-asset graph for continuity: $i=$ User Workstation, $j=$ Web/App Server, $k=$ Mid-tier Service Host, $\ell=$ Domain Controller. Binary reachability would mark both $i{\to}j{\to}\ell$ (two hops) and any direct admin path as equally “reachable.” Our probabilistic model—via $P_{2}$ (Fig. 7)—quantifies the asymmetry: high-privilege remote-admin edges dominate over multi-hop, low-privilege routes, which explains the uplift in BRE2 (and LMS2) and prioritizes the single most important control.
@@ 7.2 Illustrative Field Example: Service-Class Contrast in an Enterprise Snapshot
We complement our formal results with an illustrative observation from a production environment described by Tyagi [26]. The analysis compares adjacency induced by two services: RDP (TCP/3389), an interactive remote-admin channel, and Epic Cache (TCP/6966), an application-specific service. The induced RDP adjacency was reported as nearly complete (most node pairs reachable directly), whereas the Epic Cache adjacency was nearly empty (few pairs reachable). Under our framework, this service-class asymmetry implies substantially higher step-success and union mass for RDP-driven paths, elevating both LMSK (off-diagonal mean) and BREK (mean including the diagonal) relative to app-only services. The same field note underscores that adversaries frequently achieve lateral movement using valid accounts over remote services (RDP/SSH/SMB/WinRM) without relying on unpatched CVEs, aligning with ATT&CK’s “Remote Services” technique and our CVE-agnostic design. We emphasize that this is a single-environment qualitative example (not a statistical study). Nevertheless, the observation is consistent with our model’s predictions and with practitioner guidance prioritizing constraints on interactive remote services. [26, 2, 6, 3]
8.  Conclusion and Future Work
Conclusion.
We presented a probabilistic, service-aware model for lateral-movement exposure that cleanly separates structure (adjacency $A$) from step success (pivot potential $W$), composes hops via $\otimes$, and aggregates paths via $\oplus$. The running example and field snapshot illustrate how the model prioritizes high-impact controls beyond binary reachability.
Future work.
  • Technique/dependency modeling: relax noisy-OR independence by conditioning on attack techniques and shared prerequisites (credentials, agent presence); explore dependency-aware unions.
  • Modern environments: extend to containers, APIs, and service-mesh edges (ephemeral topologies, short-lived identities).
  • Vuln/RCE coupling: optional integration of vulnerability evidence (e.g., exploitable RCEs) as multipliers on $W$ while keeping the baseline CVE-agnostic design.
  • Business value weighting: incorporate asset criticality and blast impact (risk == probability $\times$ loss), with weighted BRE/LMS for decision support.
  • Policy optimization: compute minimal-control sets (segmentation/identity constraints) that maximally reduce $\mathrm{BRE}_{K}$ under operational constraints.